OpenStudy (zyberg):

Find the solution of this equation: x^3 + 6x = 3*sqrt(3) * x^2

Directrix (directrix):

What are the instructions for this problem?

Directrix (directrix):

If you are asked to find the roots, see if the expression will factor over the set of irrational numbers.

OpenStudy (zyberg):

Yes, @Directrix , I am asked to find the roots.

OpenStudy (eliesaab):

I will give you a hint to start doing it. Is x=0 a root?

OpenStudy (eliesaab):

You can write it as \[ x^3+ 3 x= 3 \sqrt 3 x^2\\ x^3+ 3 x-3 \sqrt 3 x^2=0\\ x( x^2 +3 -3 \sqrt 3 x)=0 \] You can finish it now

OpenStudy (eliesaab):

Use the quadratic formula for the quadratic

Directrix (directrix):

@eliesaab Do you see my error? When I factored out the x, I got: x^3 + 6x = 3*sqrt(3) * x^2 = x*( x² - ( 3√3 ) *x + 6 ) ------- ( x² - ( 3√3 ) *x + 6 ) will factor over the set of irrationals.

OpenStudy (eliesaab):

I copied the 6x as 3 x, my mistake \[ x^3+ 6 x= 3 \sqrt 3 x^2\\ x^3+ 6 x-3 \sqrt 3 x^2=0\\ x( x^2 +6 -3 \sqrt 3 x)=0 \]

OpenStudy (eliesaab):

\[ x^2-3 \sqrt{3} x+6=\left(x-2 \sqrt{3}\right) \left(x-\sqrt{3}\right) \]